OpenStudy (anonymous):

Two questions? 18.Find (fog) if f(x) = 2x - 1 and g(x) = x + 2. A. 2x - 3 B. 2x + 1 C. 2x - 1 D. 2x + 3 I think it's B (2x + 1) what I did to get my answer? I combined like terms (2x and x) = 2x or 3x? and then I added -1 + 2 and got 1. 19. What is the change that occurs to the parent function f(x) = x2 given the function f(x) = (x + 1)2 - 5. A. The graph moves 1 unit to the left and 5 units up. B. The graph moves 1 unit to the left and 5 units down. C. The graph moves 1 unit to the right and 5 units down. D. The graph moves 1 unit to the right and 5 u

OpenStudy (shaik0124):

2x+3 is answer

OpenStudy (anonymous):

The graph moves 1 unit to the right and 5 units down.

OpenStudy (shaik0124):

(fog)(x)=f(g(x))=f(x+2)=2(x+2)-1=2x+3

OpenStudy (shaik0124):

a is answer